# Unleashing the Expressive Power of Pulse-Based Quantum Neural Networks

### v1

@@ Abstract
Quantum machine learning (QML) based on Noisy Intermediate-Scale Quantum (NISQ) devices requires the optimal utilization of limited quantum resources. The commonly used gate-based QML models are convenient for software engineers, but their expressivity is restricted by the permissible circuit depth within a finite coherence time. In contrast, pulse-based models enable the construction of ”infinitely” deep quantum neural networks within the same coherence time, which may unleash greater expressive power for complex learning tasks. In this paper, we investigate this potential from the perspective of quantum control theory. We first indicate that the nonlinearity of pulse-based models comes from the encoding process that can be viewed as the continuous limit of data-reuploading in gate-based models. Subsequently, we prove that the pulse-based model can approximate arbitrary nonlinear functions when the underlying physical system is ensemble controllable. Under this condition, numerical simulations show that the expressivity can be enhanced by either increasing the pulse length or the number of qubits. As anticipated, we demonstrate through numerical examples that the pulse-based model can unleash more expressive power compared to the gate-based model. These findings establish a theoretical foundation for understanding and designing expressive QML models using NISQ devices.
I Introduction
@@
From a hardware perspective, QML models may also be parameterized by continuous-time control pulses instead of discrete-time gate parameters. Compared with gate-based models whose circuit ansatz structure must be preselected at the abstraction level, pulse-based models is natuarally this merit makes it useful for the expressivity analysis when considering the impact of the actual hardware topology.
Owing to the proximity to the underlying physical system, pulse-based models are much easier to design and implement, and thus is friendly to hardware engineers [9, 10, 11, 12, 13]. Moreover, pulse-based QML models can potentially extract more expressive power from gate-based models because they form ”infinitely” deep QNNs within the same coherence time.
@@
To the authors’ knowledge, the universal approximation property of pulse-based QNN models has not yet been explored. In this paper, we will investigate this problem from the perspective of quantum control theory, showing that a generalized controllability is required for expressivity. The remainder of this paper is arranged as follows. Section II introduces the pulse-level model as the continuous limit of gate-based data-reuploading models. Section III introduces the ensemble controllability as a sufficient condition for the examination of expressivity of pulse-based models. In Section IV, numerical experiments are performed to demonstrate the results and the effect of system dimension and the Hamiltonian structure on the model expressivity, as well as the comparison with gate-based models. Finally, concluding remarks are made in Section VI.
II From Gate-based to Pulse-based QNN models
@@
the continuous limit leads to a pulse-based QNN model described by the following Schröodinger equation:
@@
As is illustrated in Fig. 2, this model can be physically realized by a frequency tunable superconducting qubit dicated by a DC flux bias $x$. The transition between the qubit’s logical states is manipulated by the microwave Rabi control pulse $\theta(t)=\theta_{1}(t)cos\omega t+\theta_{2}(t)sin\omega t$, where $\omega$ is the frequency of the local oscillator. Since the quadrature components $\theta_{1}(t)$ and $\theta_{2}(t)$ are usually generated in a piecewise-constant waveform, the untiary transformations generated the sub-pulses can be taken as the QNN layers. Different from the gate-based models, the number of pulse-based QNN layers can be increased by shortening the sampling period $\Delta t$ without prolonging the operation time, and thus we can in principle construct “infinitely” deep QNN models.
@@
III Ensemble controllability for the expressivity of pulse-based models
@@
|$\ket{\dot{\psi}(t;\textbf{x}^{(j)})}=-i\left[H_{0}(\textbf{x}^{(j)})+\sum_{k=1}^{p}\theta_{k}(t)H_{k}\right]\ket{{\psi}(t;\textbf{x}^{(j})},$ (7)
@@
To fit the dataset, it is required that there exists a control function $\Theta(t)$ over a time period $[0,T]$ under which
@@
for all $j=1,\cdots,N$. This can be done, among non-unique choices, if the control is able to steer the quantum states to
@@
$\mathscr{L}=span\{x_{1}^{r_{1}}\cdots x_{m}^{r_{m}},r_{1},\cdots,r_{m}\in\mathbb{N}\}\otimes su(d).$ (9)
@@
$~{}[ix_{j}D_{j},ix_{1}^{r_{1}}\cdots x_{j}^{r_{j}}\cdots x_{m}^{r_{m}}X_{\alpha\beta}]$ $=$ $ix_{1}^{r_{1}}\cdots x_{j}^{r_{j}+1}\cdots x_{m}^{r_{m}}Y_{\alpha\beta},$ (10)
$~{}[ix_{j}D_{j},ix_{1}^{r_{1}}\cdots x_{j}^{r_{j}}\cdots x_{m}^{r_{m}}Y_{\alpha\beta}]$ $=$ $ix_{1}^{r_{1}}\cdots x_{j}^{r_{j}+1}\cdots x_{m}^{r_{m}}X_{\alpha\beta},$ (11)
$~{}[ix_{1}^{r_{1}}\cdots x_{j}^{r_{j}+1}\cdots x_{m}^{r_{m}}X_{\alpha\beta},iY_{\alpha\beta}]$ $=$ $x_{1}^{r_{1}}\cdots x_{j}^{r_{j}+1}\cdots x_{m}^{r_{m}}Z_{\alpha\beta},$ (12)
@@
It should be noted that the ensemble controllability is a sufficient but not necessary condition for the expressivity of pulse-based QNN models. This condition guarantees that the model is expressive no matter which measurement observable $M$ is chosen and wherever the system state is initiated. When the system is not ensemble controllable, the model may still be expressive, but this depends on the choice of the observable and the initial state.
@@ IV Numerical experiments
under which the controlled Schrodinger equation is formally invariant:
@@
over the training dataset $\{(\textbf{x}^{(k)},y^{(k)})\}$ sampled from the target function to be fitted. We start from the single-qubit model (3). The target function is chosen as the sigmoid-like function
@@
using the single-qubit model (3). It can be seen that the loss monotonically decreases with $T$ when the sampling period $\Delta t=T/K$ is fixed, but the loss stops decreasing when $T$ is above some threshold value. This implies that ”deeper” QNNs are more expressive, and the expressivity saturates when the QNN is sufficiently deep.
@@
where the coefficients are uniformly sampled in $[-30,30]$. Figure 4(b) shows that the increase of pulse duration effectively reduces the average training loss and meanwhile narrows down the distribution, which means that the ability of model for approximating arbitrary functions increases with $T$. This clearly indicates the universal approximation property of the model when the pulse duration is sufficiently long.
In addition to duration time $T$ that characterizes the “depth” of the pulse-based QNN models, the expressive power is also dependent on the number $n$ of qubits that corresponds to its width. To illustrate this, we fit the sigmoid-like function (16) using the circularly coupled multi-qubit model (6) with up to four qubits. It can be verified that this class of models is ensemble controllable as it satisfies the criterion proposed by Theorem 2. Their training-loss curves are plotted in Fig. 5 versus the pulse duration $T$, where the sampling period is fixed as $\Delta t=0.01$.
@@
V Unleashing the expressive power from gate-based models
@@
VI Concluding remarks
To summarize, we have explored the expressive power of pulsed-based QML models from a control system point of view. It is shown that the driving control pulse plays the role of data-reuploading in the continuous-time limits, and thus contributes significant nonlinearity required for the learning. We prove that pulse-based models can approximate arbitrary nonlinear functions when the underlying quantum control system is ensemble controllable. Simulations show that the expressive power increases with the model depth and width, and the pulse-based model can release more expressive power from the gate-based models.

### v2

@@ Abstract
Quantum machine learning (QML) based on Noisy Intermediate-Scale Quantum (NISQ) devices hinges on the optimal utilization of limited quantum resources. While gate-based QML models are user-friendly for software engineers, their expressivity is restricted by the permissible circuit depth within a finite coherence time. In contrast, pulse-based models enable the construction of "infinitely" deep quantum neural networks within the same time, which may unleash greater expressive power for complex learning tasks. In this paper, this potential is investigated from the perspective of quantum control theory. We first indicate that the nonlinearity of pulse-based models comes from the encoding process that can be viewed as the continuous limit of data-reuploading in gate-based models. Subsequently, we prove that the pulse-based model can approximate arbitrary nonlinear functions when the underlying physical system is ensemble controllable. Under this condition, numerical simulations demonstrate the enhanced expressivity by either increasing the pulse length or the number of qubits. As anticipated, we show through numerical examples that the pulse-based model can unleash more expressive power compared to the gate-based model. These findings lay a theoretical foundation for understanding and designing expressive QML models using NISQ devices.
1 Introduction
@@
From a hardware perspective, QML models may also be parameterized by continuous-time control pulses instead of discrete-time gate parameters. Compared with gate-based models whose circuit ansatz structure may not be hardware efficient, pulse-based models is natuarally compatible with the hardware topology. This merit makes it useful for the expressivity analysis when considering the impact of the actual hardware topology.
Owing to the proximity to the underlying physical system, pulse-based models are much easier to design and implement, and thus is friendly to hardware engineers [9, 10, 11, 12, 13]. Moreover, pulse-based QML models are potentially more expressive than the gate-based models because they form "infinitely" deep QNNs within the same coherence time.
@@
To the authors’ knowledge, the universal approximation property of pulse-based QNN models has not been explored. In this paper, we will investigate this problem from the perspective of quantum control theory, demonstrating that the pulse-based model can approximate arbitrary nonlinear functions when the underlying physical system is ensemble controllable. Numerical simulations show that the pulse-based model can unleash more expressive power from the gate-based model.
The remainder of this paper is arranged as follows. Section 2 introduces the pulse-level model as the continuous limit of gate-based data-reuploading models. Section 3 introduces the ensemble controllability as a sufficient condition for the examination of expressivity of pulse-based models. In Section 4, numerical experiments are performed to demonstrate the results and the effect of system dimension and the Hamiltonian structure on the model expressivity, as well as the comparison with gate-based models. Finally, concluding remarks are made in Section 5.
2 From gate-based to pulse-based QNN models
@@
the continuous limit leads to a pulse-based QNN model described by the following Schrödinger equation:
@@
As is illustrated in Fig. 2, this model can be physically realized by a frequency tunable superconducting qubit dicated by a DC flux bias $x$. The transition between the qubit’s logical states is manipulated by the microwave Rabi control pulse $\theta(t)=\theta_{1}(t)\cos\omega t+\theta_{2}(t)\sin\omega t$, where $\omega$ is the frequency of the local oscillator. Since the quadrature components $\theta_{1}(t)$ and $\theta_{2}(t)$ are usually generated in a piecewise-constant waveform, the unitary transformations generated by the sub-pulses can be taken as the QNN layers. Different from the gate-based models, the number of pulse-based QNN layers can be increased by shortening the sampling period $\Delta t$ without increasing the operation time, and thus we can in principle construct "infinitely" deep QNN models.
@@
3 Ensemble controllability for the expressivity of pulse-based models
@@
|$\ket{\dot{\psi}(t;\textbf{x}^{(j)})}=-i\left[H_{0}(\textbf{x}^{(j)})+\sum_{k=1}^{p}\theta_{k}(t)H_{k}\right]\ket{{\psi}(t;\textbf{x}^{(j})},$
@@
To perfectly fit the dataset, there should exist a control function $\Theta(t)$ over a time period $[0,T]$ under which
@@
for all $j=1,\cdots,N$. This can be done if the control is able to steer the quantum states to
@@
$\mathscr{L}=span\{x_{1}^{r_{1}}\cdots x_{m}^{r_{m}},r_{1},\cdots,r_{m}\in\mathbb{N}\}\otimes su(d).$
@@
$\leavevmode\nobreak\ [ix_{j}D_{j},ix_{1}^{r_{1}}\cdots x_{j}^{r_{j}}\cdots x_{m}^{r_{m}}X_{\alpha\beta}]$ $=$ $ix_{1}^{r_{1}}\cdots x_{j}^{r_{j}+1}\cdots x_{m}^{r_{m}}Y_{\alpha\beta},$ (8)
$\leavevmode\nobreak\ [ix_{j}D_{j},ix_{1}^{r_{1}}\cdots x_{j}^{r_{j}}\cdots x_{m}^{r_{m}}Y_{\alpha\beta}]$ $=$ $ix_{1}^{r_{1}}\cdots x_{j}^{r_{j}+1}\cdots x_{m}^{r_{m}}X_{\alpha\beta},$ (9)
$\leavevmode\nobreak\ [ix_{1}^{r_{1}}\cdots x_{j}^{r_{j}+1}\cdots x_{m}^{r_{m}}X_{\alpha\beta},iY_{\alpha\beta}]$ $=$ $x_{1}^{r_{1}}\cdots x_{j}^{r_{j}+1}\cdots x_{m}^{r_{m}}Z_{\alpha\beta},$ (10)
@@
It should be noted that the ensemble controllability is a sufficient but not necessary condition for the expressivity of pulse-based QNN models. This condition guarantees that the model is expressive no matter which measurement observable $M$ is chosen and no matter what initial state is selected. When the system is not ensemble controllable, the model may still be expressive, but this depends on the choice of the proper observable and the initial state.
@@ 4 Numerical experiments
under which the controlled Schrödinger equation is formally invariant:
@@
over the training dataset $\{(\textbf{x}^{(k)},y^{(k)})\}$ sampled from the target function to be fitted.
4.1 The approximation of nonlinear functions
We start from the single-qubit model (3). The target function is chosen as the sigmoid-like function
@@ 4.2 The affection of pulse duration
using the single-qubit model (3). It can be seen that the loss monotonically decreases with $T$ when the sampling period $\Delta t=T/K$ is fixed, but the loss stops decreasing when $T$ is above some threshold value. This implies that "deeper" QNNs are more expressive, and the expressivity saturates when the QNN is sufficiently deep.
@@
where the coefficients are uniformly sampled in $[-30,30]$. Figure 4(b) shows that the increase of pulse duration effectively reduces the average training loss and meanwhile narrows down the distribution, which means that the ability of the model to approximate arbitrary functions increases with $T$. This clearly indicates the universal approximation property of the model when the pulse duration is sufficiently long.
In addition to the duration time $T$ that characterizes the "depth" of the pulse-based QNN models, the expressive power is also dependent on the number $n$ of qubits that corresponds to its width. For illustration, we fit the sigmoid-like function (14) using the circularly coupled multi-qubit model (6) with up to four qubits. It can be verified that this class of models is ensemble controllable as it satisfies the criterion proposed by Theorem 2. Their training-loss curves are plotted in Fig. 5 versus the pulse duration $T$, where the sampling period is fixed as $\Delta t=0.01$.
@@
4.3 Unleashing the expressive power from gate-based models
@@
5 Concluding remarks
To summarize, we have explored the expressive power of pulse-based QML models from a control system point of view. It is shown that the driving control pulse plays the role of data-reuploading in the continuous-time limit, and thus contributes significant nonlinearity required for the learning. We prove that pulse-based models can approximate arbitrary nonlinear functions when the underlying quantum control system is ensemble controllable. Simulations show that the expressive power increases with the model depth and width, and the pulse-based model can release more expressive power from the gate-based models.